OpenStudy (anonymous):

please help me with this question, no answers allowed! :) Is the simplified form of 2square root of 3 − 2square root of 3 rational? Yes No

OpenStudy (anonymous):

NOOOO

zepdrix (zepdrix):

This is the problem?\[\Large\rm 2\sqrt3-2\sqrt3\]What numerical value does it equal? Remember how to combine like-terms? :)

OpenStudy (anonymous):

so yes, it is IRRATIONAL, concluding the answer as NO? @zepdrix

zepdrix (zepdrix):

Think of the sqrt3 as just like uhhh... potato. So your problem is really 2potato minus 2potato. Which ... equals zero, ya? Is zero a rational number?

zepdrix (zepdrix):

Rational numbers are numbers which we can write as a ratio of integers. Examples: 5 is a rational number because we can write it as \(\large\rm \frac{5}{1}\). whole number divided by another whole number. 2.5 is a rational number because we can write it as \(\large\rm \frac{5}{2}\). 0.001 is a rational number because we can write it as \(\large\rm \frac{1}{1000}\).

zepdrix (zepdrix):

But what about zero? :)

OpenStudy (anonymous):

wait, but the sqr root of 3 is a decimal, concluding that it is IRRATIONAL... I am just solving it, and not sure where you get the numbers lol

zepdrix (zepdrix):

You need to simplify BEFORE you determine whether or not it's rational.

zepdrix (zepdrix):

\[\large\rm 2\sqrt3-2\sqrt3=0\]When we do the subtraction, we get zero. We have simplified the expression. We no longer care about the fact that the sqrt3's were irrational.

zepdrix (zepdrix):

Only look at the zero.

zepdrix (zepdrix):

So... determine whether or not zero is a rational number, and you will have your answer.